OpenStudy (anonymous):

Given \sin(\alpha)=\frac{4}{7} and \alpha is in quadrant II

OpenStudy (anonymous):

\[\Huge \sin(\alpha)=\frac{4}{7}\]

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok well do you know to do the \[\Large \sin^{-1} ~function~on~your~calculator?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well do that and put 4/7 and get your answer and if you know the acrobym astc then you know that this value will be positive

OpenStudy (anonymous):

not sure what acrobym astc is.. sorry

OpenStudy (anonymous):

It just tells you what value will be positive in each quadrant. 

OpenStudy (anonymous):

And astc can be remembered by "All Students Take Calculus"